# MUSE-FM: Multi-task Environment-aware Foundation Model for Wireless Communications

### v1

@@ III-D2 Post-processor and unified data decoder
Finally, the post-processor extracts the corresponding output for the target task:
@@ III-E Foundation Model Backbone
In this subsection, we elaborate on the architecture of the transformer-based backbone. Position embeddings are added to the combined tokens $\mathrm{X}_{\rm FM}^{\rm emb}$, providing sequential information. Thus, the input to the transformer blocks is denoted as $\mathrm{X}_{\rm FM}^{I(1)}$. As described in the scene encoder, each transformer block has two sub-layers. The first is a multi-head self-attention mechanism, and the second is a simple MLP network. The $l$-th transformer block takes the output of the l1l-1-th transformer block as input, that is, $\mathrm{X}_{\rm FM}^{I(l)}=\mathrm{X}_{\rm FM}^{O(l-1)}$. LN operates independently across the feature dimension for each token in the sequence and each sample in the batch:
@@ III-F Loss Function and Training Schedule
where ${\rm loss}_{n}$ is the loss function of task $n$. Besides, due to the disparities in the convergence rates and loss scales across tasks, we implement a weighted linear combination scheme with $\alpha_{n}$. Next, we briefly illustrate the loss function for each task. For multi-user precoding, we directly optimize the negative sum rate in an unsupervised manner, which is presented in (9). For regression problems including channel estimation, user localization, and MIMO detection, the MSE loss is selected as the loss function. For channel decoding, following [20], we utilize the cross entropy loss to optimize the estimated binary multiplicative noise. The selected tasks, the corresponding training metrics and evaluation metrics are summarized in Table I.
@@
where $\beta_{n}$ and ${\rm loss}_{val,n}$ denote the weight and loss of the task $n$, respectively. For the validation loss, NMSE is utilized to evaluate the performance of channel estimation and MIMO detection; while the normalized average positioning error is chosen to provide intuitive evaluation for user localization. For multi-user precoding, the loss function is the negative sum rate; for channel decoding, the loss function is the bit error ratio (BER).
@@ IV Simulation Results
In this section, extensive numerical simulations are presented to verify the effectiveness of the proposed method. First, we introduce the constructed multi-task multi-scenario multi-modal dataset. The detailed simulation setup and training configurations are then illustrated. Finally, we provide thorough performance evaluation and analysis of the proposed MUSE-FM.
@@ IV-A2 Multi-task dataset
Sionna [24], an open-source library for link-level simulations based on TensorFlow, is employed to obtain datasets for multiple tasks in constructed environments. In particular, the BS is placed at (4.75m,4.75m,2.5m)(-4.75m,4.75m,2.5m), which remains fixed in all scenes, while the users are randomly distributed in the scenario. Then the Sionna ray tracing simulator [25] is implemented to generate multipath channels in 3D environments, accurately representing propagation conditions by accounting for reflection, diffraction, and scattering. The simulated paths between the BS and the UE are also illustrated in Fig. 5. The parameters of the system are listed in Table II. Based on the generated channels, we can calculate the data samples for different tasks in the MIMO-OFDM transceiver, including the received signal, user positions, multi-user channels, etc.
@@ IV-B Simulation Setup and Training Details
For hyperparameters in network training, we set the batch size as 100, and the initialized learning rate as 0.0001 with a cosine decay scheduler. We utilize Adam optimizer for model training with $\rm betas=(0.9,0.999)$. The model undergoes training for 500 epochs over the multi-task dataset. The smallest version of GPT-2 [4] with feature dimension $D=768$ is adopted as the backbone, where all parameters are trained to obtain a foundation model dedicated for wireless communications. All training and inference of the proposed model is conducted on four NVIDIA GeForce RTX 4090 24GB GPUs. To validate the superiority of the proposed method, several baselines are implemented which can be categorized into the following groups.
Single-task small models. This class of methods employ conventional DL architectures, such as MLPs, CNNs, and transformers, to address specific downstream tasks, which often have a relatively small number of parameters.
@@
Single-task FMs. These methods typically finetune FMs for a single downstream task, which leverage the powerful modeling capabilities of FMs to improve the performance.
@@ IV-C3 Scaling behavior
We summarize the main insights for the experiments in this subsection.
@@ IV-C4 Impact of prompt-guided unified data encoder-decoder pair
In this subsection, the effectiveness of the proposed prompt-guided unified data encoder-decoder pair is evaluated. As discussed in Section III-C, designing separate data encoders and decoders as [17] for each task and parameter helps task alignment and facilitates multi-task learning, but it incurs high computational and deployment overhead, given the diverse tasks and parameter variations. Besides, when encountering a new task or parameter configuration with different input/output formats, FMs with task-specific data encoder-decoder pairs may not work. Instead, a new task-specific encoder-decoder pair is required to be designed and trained, which limits the scalability of the models. Therefore, we propose a unified prompt-guided data encoder-decoder pair to improve the scalability of multi-task FM, while retaining superior performance for all tasks. To validate the effectiveness, we first compare the proposed MUSE-FM with multi-task FM with task-specific encoder-decoder pairs in [17]. Then we evaluate the generalization ability of the unified prompt-guided data encoder-decoder pair for a new parameter configuration.
@@
Fig. 7 utilizes channel estimation for instance to illustrate the ability of the proposed MUSE-FM in handling new task configurations. During the training process, the pilot length is set as 4, while the pilot length is changed to 6 during evaluation. Varying pilot quantities alter input dimensions, rendering FM with task-specific encoder-decoder pairs inapplicable. In contrast, our proposed method remains directly applicable to new parameter configurations. To comprehensively demonstrate its capability, except the existing baselines, we compare: (1) FM trained specifically for pilot length=6, (2) Zero-shot performance where FM trained with pilot length=4 is applied to pilot length=6, and (3) Few-shot performance where FM trained with pilot length=4 is fine-tuned with 1% data for 5 epochs. As shown in Fig. 7, the zero-shot performance of our proposed method surpasses ReEsNet and LS-based methods and achieves comparable performance with CE-ViT. With only minimal fine-tuning (1% data for 5 epochs), our method outperforms all baselines, indicating the scalability of the proposed unified prompt-guided encoder-decoder pair.
In conclusion, we observe that the proposed prompt-guided unified data encoder-decoder pair achieves comparable or better performance than task-specific encoder-decoder pairs, while improves the adaptability of the multi-task FM.
@@ V Conclusions
In this work, we propose a multi-task environment-aware FM, namely MUSE-FM for wireless communications with a unified architecture to handle multiple tasks and incorporate scenario information. Through multi-modal input integration of environmental context and wireless data, the proposed model manages to facilitate cross-scenario feature extraction and enhances the environmental adaptability. The proposed prompt-guided data encoder-decoder pair employs instructions to dynamically generate task-specific parameters, and thus handles data with heterogeneous formats and distributions across multiple tasks with a unified encoder-decoder pair. Simulation results have demonstrated the multi-task learning and generalization capabilities of the proposed method. Besides, incorporating environmental information facilitates cross-scenario learning ability, while the proposed prompt-guided unified encoder-decoder pair improves the scalability of the model. Furthermore, with limited data samples, MUSE-FM harnesses cross-task dependencies for knowledge transfer and enhance joint optimization compared with task-specific FMs. Future works can focus on more effective use of environmental information. In addition, the end-to-end optimization for the transceiver can be employed to further improve the performance.

### v2

@@ III-D2 Post-processor and unified data decoder
Finally, the post-processor extracts the corresponding output for the target task, which can be expressed as
@@ III-E Foundation Model Backbone
The FM backbone fuses multi-modal inputs and effectively extracts features for multi-task learning. Leveraging the Transformer’s attention mechanism, the FM dynamically assigns varying importance to tokens representing different environmental features. This enables it to distill task-beneficial representations from a unified prior knowledge of channel distribution $\mathrm{X}_{s}^{\rm emb}$ for multiple tasks.
Architecturally, it is composed of stacked Transformer modules, which is elaborated in detail. Position embeddings are added to the combined tokens $\mathrm{X}_{\rm FM}^{\rm emb}$, providing sequential information. Thus, the input to the transformer blocks is denoted as $\mathrm{X}_{\rm FM}^{I(1)}$. As described in the scene encoder, each transformer block has two sub-layers. The first is a multi-head self-attention mechanism, and the second is a simple MLP network. The $l$-th transformer block takes the output of the l1l-1-th transformer block as input, that is, $\mathrm{X}_{\rm FM}^{I(l)}=\mathrm{X}_{\rm FM}^{O(l-1)}$. LN operates independently across the feature dimension for each token in the sequence and each sample in the batch:
@@ III-F Loss Function and Training Schedule
where ${\rm loss}_{n}$ is the loss function of task $n$. Besides, due to the disparities in the convergence rates and loss scales across tasks, we implement a weighted linear combination scheme with $\alpha_{n}$. Specifically, $\alpha_{n}$ are determined with Dynamic Weight Average (DWA) algorithm [24], to dynamically adjust each task’s weight based on its loss every epoch. Next, we briefly illustrate the loss function for each task. For multi-user precoding, we directly optimize the negative sum rate in an unsupervised manner, which is presented in (9). For regression problems including channel estimation, user localization, and MIMO detection, the MSE loss is selected as the loss function. For channel decoding, following [19], we utilize the cross entropy loss to optimize the estimated binary multiplicative noise. The selected tasks, the corresponding training metrics and evaluation metrics are summarized in Table I.
@@
where $\beta_{n}$ and ${\rm loss}_{val,n}$ denote the weight and loss of the task $n$, respectively. Specifically, NMSE is utilized to evaluate the performance of channel estimation and MIMO detection; while the normalized average positioning error is chosen to provide intuitive evaluation for user localization. For multi-user precoding, the loss function is the negative sum rate; for channel decoding, the loss function is the bit error ratio (BER).
@@ IV Simulation Results
In this section, extensive numerical simulations are presented to verify the effectiveness of the proposed method. First, we introduce the constructed multi-task multi-scenario multi-modal dataset. The detailed simulation setup and training configurations are then illustrated. Finally, we provide thorough performance evaluation and analysis of MUSE-FM.
@@ IV-A2 Multi-task dataset
Sionna [26], an open-source library for link-level simulations based on TensorFlow, is employed to obtain datasets for multiple tasks in constructed environments. In particular, the BS is randomly placed at $(\pm 4.75m,\pm 4.75m,2.5m)$, and the element in the scene graph matrix is set as 2 to label the BS position. Besides, the users are randomly distributed in the scenario. Then the Sionna ray tracing simulator [27] is implemented to generate multipath channels in 3D environments, accurately representing propagation conditions by accounting for reflection, diffraction, and scattering. The simulated paths between the BS and the UE are also illustrated in Fig. 5. The parameters of the system are listed in Table II. Based on the generated channels, we can calculate the data samples for different tasks in the MIMO-OFDM transceiver, including the received signal, user positions, channels, etc.
@@ IV-B Simulation Setup and Training Details
For hyperparameters in network training, we set the batch size as 100, and the initialized learning rate as 0.0001 with a cosine decay scheduler. We utilize Adam optimizer for model training with $\rm betas=(0.9,0.999)$. The model undergoes training for 500 epochs over the multi-task dataset. The smallest version of GPT-2 [4] with feature dimension $D=768$ is adopted as the backbone, where all parameters are trained to obtain a FM dedicated for wireless communications. All training and inference of the proposed model is conducted on four NVIDIA GeForce RTX 4090 24GB GPUs. To validate the superiority of the proposed method, several baselines are implemented which can be categorized into the following groups. To enhance the model’s ability to process scene graphs of diverse sizes and resolutions, we apply data augmentation by resizing scene graphs to multiple scales during training.
Single-task small models. This class of methods employ conventional DL architectures, such as MLPs, CNNs, and transformers, to address different tasks individually, which often have a relatively small number of parameters.
@@
Single-task FMs. These methods typically finetune FMs for a single task, which leverage the powerful modeling capabilities of FMs to improve the performance.
@@ IV-C2 Task-specific performance analysis
Moreover, MUSE-FM demonstrates advantages in training and inference overhead compared to task-specific FMs. Regarding memory overhead, with 134.44M parameters, MUSE-FM incurs significantly lower storage costs compared to deploying five task-specific FMs (each 93.20M parameters, totaling 466M parameters). Compared to the sequential training and inference of multiple task-specific FMs, MUSE-FM reduces training latency by 44.3% and inference latency by 48.9%.
@@ IV-C3 Scaling behavior
We summarize the main insights for the experiments in this subsection as follows.
@@ IV-C4 Impact of prompt-guided unified data encoder-decoder pair
In this subsection, the effectiveness of the proposed prompt-guided unified data encoder-decoder pair is evaluated. As discussed in Section III-C, designing separate data encoders and decoders as [17] for each task and parameter helps task alignment and facilitates multi-task learning, but it incurs high computational and deployment overhead, given the diverse tasks and parameter variations. Besides, when encountering a new task or parameter configuration with different input/output formats, FMs with task-specific data encoder-decoder pairs may not work. Instead, a new task-specific encoder-decoder pair is required to be designed and trained, which limits the scalability of the models. Therefore, we propose a unified prompt-guided data encoder-decoder pair to improve the scalability of multi-task FM, while retaining superior performance for all tasks. To validate the effectiveness, we compare the proposed MUSE-FM with multi-task FM with task-specific encoder-decoder pairs in [17].
@@
In conclusion, we observe that the proposed prompt-guided unified data encoder-decoder pair achieves comparable performance with task-specific encoder-decoder pairs, while improves the adaptability of the multi-task FM.
IV-C5 Generalization ability
Then we evaluate the generalization ability of MUSE-FM for a new parameter configuration (e.g. antenna number, pilot length, user number). Fig. 7 utilizes channel estimation for instance to illustrate the ability of the proposed MUSE-FM in handling new task configurations. During the training process, the pilot length is set as 4, while the pilot length is changed to 6 during evaluation. Varying pilot quantities alter input dimensions, rendering FM with task-specific encoder-decoder pairs inapplicable. In contrast, our proposed method remains directly applicable to new parameter configurations. To comprehensively demonstrate its capability, except the existing baselines, we compare: (1) FM trained specifically for pilot length = 6, (2) Zero-shot performance where FM trained with pilot length = 4 is applied to pilot length = 6, and (3) Few-shot performance where FM trained with pilot length = 4 is fine-tuned with 1% data for 5 epochs. As shown in Fig. 7, the zero-shot performance of our proposed method surpasses ReEsNet and LS-based methods and achieves comparable performance with CE-ViT. With only minimal fine-tuning (1% data for 5 epochs), our method outperforms all baselines, indicating the generalization ability of the proposed MUSE-FM .
@@ V Conclusions
In this work, we propose a multi-task environment-aware FM, namely MUSE-FM for wireless communications with a unified architecture to handle multiple tasks and incorporate scenario information. Through multi-modal input integration of environmental context and wireless data, the proposed model manages to facilitate cross-scenario feature extraction and enhances the environmental adaptability. The proposed prompt-guided data encoder-decoder pair employs instructions to dynamically generate task-specific parameters, and thus handles data with heterogeneous formats and distributions across multiple tasks with a unified encoder-decoder pair. Simulation results have demonstrated the multi-task learning and generalization capabilities of the proposed method. Besides, incorporating environmental information facilitates cross-scenario learning ability, while the proposed prompt-guided unified encoder-decoder pair improves the scalability of the model. Furthermore, with limited data samples, MUSE-FM harnesses cross-task dependencies for knowledge transfer and enhance joint optimization compared with task-specific FMs. Future works can focus on more effective use of environmental information. In addition, the end-to-end optimization for the transceiver can be used to further improve the performance.